OpenStudy (anonymous):

2. Find the area of the given circle. Round to the nearest tenth. (1 point) 11.0 ft2 38.5 ft2 42.8 ft2 113.0 ft2

OpenStudy (anonymous):

pic.

OpenStudy (anonymous):

all i need to know is that 38.5 ft2. yes or no!!!!!!!!!

OpenStudy (raden):

no, recalculate again... area = pi * r^2 = 3.14 * 6^2 = ...

OpenStudy (anonymous):

no i dont think that is right i believe you find it out by doing 12 times pi